Bill is 3 years older than Carlos. Five years ago, Bill was twice Carlos's age. How old are they now? Write a system to represent this problem.
Answer by lwsshak3(11628)

Bill is 3 years older than Carlos. Five years ago, Bill was twice Carlos's age. How old are they now? Write a system to represent this problem.
***
Now:
x=Carlos'age
x+3=Bill's age
..
5 yrs ago:
x-5=Carlos' age
x+3-5=x-2=Bills age
x-2=2(x-5)
x-2=2x-10
x=8
x+3=11
Carlos'age now=8
Bill's age now=11